# LLINBO: Trustworthy LLM-in-the-Loop Bayesian Optimization

### v1

@@ A.1 Proof of Theorem 1
where $\mathcal{F}_{t-1}$ denotes the filtration until $t-1$ and $\nu_{t}=\mathbb{E}[r_{t}|\mathcal{F}_{t-1},B_{t}]$.
@@ A.2 Proof of Theorem 2
By assuming that $\psi_{t}=\mathcal{O}(1/\sqrt{t})$ and by the Lemma 4 in [8], which allows us to bound the sum of variance at the evaluated designs, we have
@@ A.3 Proof of Theorems 3 and 4
Lemma 6 (Appendix C in [8]).
@@
where $K_{t-1}$ denotes the Gram matrix at time $t$, defined identically as in the main paper but indexed with a subscript to emphasize its dependence on the data available up to time $t-1$. Next, we derive the AF via models constructed by $\mathcal{D}_{t-1}\cup\{(x_{\text{LLM},t},\tilde{f}_{t-1,s}(x_{\text{LLM},t}))\}$, which we denoted those models as $\mathcal{M}_{t,s}\ \forall s\in I_{t}$.
@@
(Lemma 1 in [6]) Assuming $\mathbb{E}_{\mathcal{M}_{t,s}}[\alpha(x,\mathcal{M}_{t,s})]$ exists, and there exists a function $a:\mathbb{R}\rightarrow\mathbb{R}$ such that
@@
$\ \ \ \ \ +|k^{+}_{t-1}(x)^{\top}(K^{+}_{t-1}+\hat{\lambda}_{t}I)^{-1}[\delta_{1},...,\delta_{t-1},\tilde{\delta}]^{\top}|\hfill$ (by (4))
@@
$\left|{k^{+}_{t-1}(x)}^{\top}(K^{+}_{t-1}+\hat{\lambda}_{t}I)^{-1}[0,...,0,\tilde{\delta}]^{\top}\right|$ $\leq\hat{\lambda}_{t}^{-1/2}\,\sigma^{+}_{t-1}(x)\,\sqrt{\begin{bmatrix}0&\tilde{\delta}\end{bmatrix}K^{+}_{t-1}(K^{+}_{t-1}+\hat{\lambda_{t}}I)^{-1}\begin{bmatrix}0&\tilde{\delta}\end{bmatrix}^{\top}}.$
@@
Appendix B Numerical Experiments Details
We utilize GPT-3.5-turbo as the LLM agent, selected for its demonstrated capability to generate high-quality responses. The temperature parameter is set to its default value of 1.0. Prompt structures for LLAMBO are primarily adapted from the methodology proposed by [25]. For each task, we define a task-specific system prompt. Specifically, the system prompt for black-box optimization is: "You are an AI assistant that helps people find the maximum of a black-box function." and for hyperparameter tuning tasks: "You are an AI assistant that helps me reduce the mean square error by tuning the hyperparameters in a machine learning model."
We use SingleTaskGP in Python’s BOTorch package [3] as the surrogate model when a statistical model is involved. Namely, its prior mean is set to be constant, where the constant is learned while training, and the kernel function is set to be martern-52 with automatic relevance determination.
B.1 Experimental Details for Black-box Optimization Task
@@
Prompts design for black-box optimization task.
@@
The Function Patterns included in each Description Card are derived from the benchmark function descriptions provided by [33], and a summary of these patterns is presented in Table 1.
@@
[…]
@@
In the LLAMBO framework, candidate sampling is facilitated by a structured prompt designed to elicit a diverse set of potential query points. This mechanism is illustrated in the Candidate sampling phase of Table 2. At each iteration, we prompt LLM 10 times to generate a total of 10 candidate points. To enhance the diversity of these candidates, we follow the strategy outlined in [25], where the content of the Data Card is permuted across prompts.
The LLAMBO framework [25] introduces a hyperparameter $\alpha=0.1$ to balance exploration and exploitation during the candidate sampling phase. At iteration $t$, we compute the Target Score based on the current observed values $\{y_{i}\}$ as follows:
@@
In the LLAMBO framework, a surrogate prompt is used to estimate the predictive mean and variance at each candidate point generated by the candidate sampling prompt. This process corresponds to the Surrogate modeling phase illustrated in Table 2. To promote variability in the surrogate responses, we similarly permute the Data Card across prompts. Finally, an AF is applied to select the next query point. We adopt the Expected Improvement (EI) criterion [17], consistent with the acquisition strategy employed in [25].
@@
Table 2: Prompts used across different stages of LLAMBO and LLAMBO-light in the black-box optimization task.
B.2 Experiment Details for Hyperparameter Tuning Task
@@
[…]
  • max_depth (Maximum depth of a tree): $[-1,50]$ (integer; $-1$ indicates no limit)
@@
The prompt settings for both LLAMBO and LLAMBO-light in the hyperparameter tuning task follow the same configuration as in the black-box optimization task ($\alpha$ and AF), with the exception of the prompt structure. In particular, the hyperparameter tuning prompts also require both the Description Card and the Data Card to capture the relevant model specifications and historical evaluations.
@@
Appendix C 3D Printing Details
@@
C.1 Qualifying the Stringing Percentage
@@
C.2 Prompts Design
The settings of LLMs are the same as in Appendix B.1. The system prompt is You are an AI assistant that helps me optimize the 3D manufacturing process by controlling parameters. An example of the Data Card is "(Nozzle Temperature, Z Hop Height, Coasting Volume, Retraction Distance, Outer Wall Wipe Distance): (235, 0.3, 0.06, 4, 0.3), Stringing percentage: 12%. We also need a Parameter Description Card to describe the controllable and fixed variables, which is

### v2

@@ B.1 Proof of Theorem 1
where $\mathcal{F}_{t-1}$ denotes the filtration until t1t-1 and $\nu_{t}=\mathbb{E}[r_{t}|\mathcal{F}_{t-1},B_{t}]$.
@@ B.2 Proof of Theorem 2
By assuming that $\psi_{t}=\mathcal{O}(1/\sqrt{t})$ and by the Lemma 4 in Chowdhury & Gopalan (2017), which allows us to bound the sum of variance at the evaluated designs, we have
@@ B.3 Proof of Theorems 3 and 4
Lemma 6 (Appendix C in Chowdhury & Gopalan (2017)).
@@
where $K_{t-1}$ denotes the Gram matrix at time $t$, defined identically as in the main paper but indexed with a subscript to emphasize its dependence on the data available up to time t1t-1. Next, we derive the AF via models constructed by $\mathcal{D}_{t-1}\cup\{(x_{\text{LLM},t},\tilde{f}_{t-1,s}(x_{\text{LLM},t}))\}$, which we denoted those models as $\mathcal{M}_{t,s}\ \forall s\in I_{t}$.
@@
(Lemma 1 in Chen et al. (2025)) Assuming $\mathbb{E}_{\mathcal{M}_{t,s}}[\alpha(x,\mathcal{M}_{t,s})]$ exists, and there exists a function $a:\mathbb{R}\rightarrow\mathbb{R}$ such that
@@
[…]
@@
[…]
@@
Appendix C Additional Experiments on HAIC BO
@@
(a) Branin-2DD
(b) Levy-2DD
[…]
Appendix D Selection between the Proposed Algorithms and Hyperparameters
Choosing between the proposed methods.
It is noteworthy that the regret bounds for all three methods contain no variables or assumptions on the LLMs, thereby ensuring the no-harm guarantees introduced by Xu et al. (2024). In other words, the quality of LLM suggestions does not degrade their performance, and the choice among them can therefore be guided by practical needs. LLINBO-Transient is the most interpretable and practical for non-expert users, employing an explicit temporal schedule to reduce LLM influence over time. Importantly, the reliance on LLM suggestions diminishes as the probability of querying the LLM approaches zero, making this variant suitable for practitioners prioritizing transparency, simplicity, or scenarios where accessing LLMs is costly. LLINBO-Justify adopts a more data-driven approach by learning a justification threshold for LLM suggestions without altering the BO machinery, thereby maintaining interpretability while offering adaptive control—an attractive option when flexibility is desired without structural changes. Finally, LLINBO-Constrained is the most robust and theoretically grounded variant, integrating LLMs and BO through a probabilistic constraint that automatically hedges against finite-sample uncertainty and requires no additional hyperparameter tuning, making it particularly well-suited for safety-critical or resource-constrained settings where minimizing risk and avoiding hyperparameter tuning are essential.
Selecting hyperparameters.
We acknowledge that leveraging LLMs for BO is still in its early stages. As such, tuning the algorithm’s parameters based on the LLM’s level of understanding remains an open but important research direction. Nevertheless, we outline below general-purpose strategies for selecting these parameters. For $p_{t}$ in LLINBO-Transient, our approach introduces a diminishing reliance on the LLM over time. We therefore set $1-p_{t}\in\mathcal{O}(1/t^{2})$, which limits the influence of potentially unreliable LLM suggestions as optimization progresses. Indeed, we consistently observed that the LLM’s ability to exploit diminishes rapidly over time, unsurprising since LLMs lack explicit surrogate modeling and calibrated uncertainty; however, when the problem domain is well understood by the LLM (e.g., hyperparameter tuning on standard datasets), the increase in $p_{t}$ can be made more gradual. In contrast, the performance of the LLM is less critical in LLINBO-Justify, as this variant is primarily data-driven and can automatically hedge against unreliable suggestions. Following our theoretical results, we recommend using a conservative decreasing schedule for $\psi_{t}$ ($\mathcal{O}(1/t)$, Theorem 2 in the main paper) and setting $\psi_{0}$ in a way that reflects the structure of the acquisition function. For instance, when using UCB, $\psi_{0}$ can be the posterior variance at the first LLM-suggested point, or in the case of Thompson Sampling, the difference between the maximum and minimum values in a posterior sample. Finally, LLINBO-Constrained was specifically designed to minimize the need for hyperparameter tuning, with the only parameter being the sampling size from the constrained GP, which should be dictated by available computational resources. We recommend starting with the largest feasible sample size and then gradually reducing it (as permitted by our theory) based on constraints. In our implementation, we began with a large sample size and reduced it at a rate of $\mathcal{O}(1/t^{2})$, which offered a good balance between computational efficiency and performance.
Appendix E Numerical Experiments Details
We utilize GPT-3.5-turbo as the LLM agent, selected for its demonstrated capability to generate high-quality responses. The temperature parameter is set to its default value of 1.0. Prompt structures for LLAMBO are primarily adapted from the methodology proposed by Liu et al. (2024). For each task, we define a task-specific system prompt. Specifically, the system prompt for BBO is: "You are an AI assistant that helps people find the maximum of a black-box function." and for hyperparameter tuning tasks: "You are an AI assistant that helps me reduce the mean square error by tuning the hyperparameters in a machine learning model."
We use SingleTaskGP in Python’s BOTorch package Balandat et al. (2020) as the surrogate model when a statistical model is involved. Namely, its prior mean is set to be constant, where the constant is learned while training, and the kernel function is set to be matern 5/2 with automatic relevance determination.
E.1 Experimental Details for BBO
For the BBO task, we employ the following simulation functions: Levy-2DD, Rastrigin-2DD, Branin-2DD, Bukin-2DD, Hartmann-4DD, and Ackley-6DD, as implemented in the Virtual Library of Simulation Experiments Surjanovic & Bingham (2013). Each function is rescaled to the unit hypercube $0,1]^{D}$ , and a negative sign is applied to the response to convert the problem into a maximization task. A summary of these simulation functions is provided below.
  • Levy-2DD
@@ E.1 Experimental Details for BBO
Prompts design for BBO task.
@@
The Function Patterns included in each Description Card are derived from the benchmark function descriptions provided by Surjanovic & Bingham (2013), and a summary of these patterns is presented in Table 1.
@@
Levy-2DD highly multimodal but with a unique global maximum.
Rastrigin-2DD which is highly multimodal, non-convex function with a large number of regularly spaced local minima.
Branin-2DD smooth, multimodal benchmark with three global maxima
Bukin-2DD steep, narrow, and highly non-convex landscape with a sharp valley and a unique global maximum
Hartmann-4DD 4-dimensional, non-convex, multi-modal and is composed of weighted, anisotropic Gaussian-like bumps centered at different points, making it highly non-separable and challenging to optimize.
Ackley-6DD 6-dimensional, non-convex, and multi-modal. The function exhibits a nearly flat outer region and a large hole at the center, resulting in many local optima surrounding a single global optimum. It is highly symmetric and separable in nature, but optimization is still challenging due to the numerous local maxima.
@@
In the LLAMBO framework, candidate sampling is facilitated by a structured prompt designed to elicit a diverse set of potential query points. This mechanism is illustrated in the Candidate sampling phase of Table 2. At each iteration, we prompt LLM 10 times to generate a total of 10 candidate points. To enhance the diversity of these candidates, we follow the strategy outlined in Liu et al. (2024), where the content of the Data Card is permuted across prompts.
The LLAMBO framework Liu et al. (2024) introduces a hyperparameter $\alpha=0.1$ to balance exploration and exploitation during the candidate sampling phase. At iteration $t$, we compute the Target Score based on the current observed values $\{y_{i}\}$ as follows:
@@
In the LLAMBO framework, a surrogate prompt is used to estimate the predictive mean and variance at each candidate point generated by the candidate sampling prompt. This process corresponds to the Surrogate modeling phase illustrated in Table 2. To promote variability in the surrogate responses, we similarly permute the Data Card across prompts. Finally, an AF is applied to select the next query point. We adopt the Expected Improvement (EI) criterion Jones et al. (1998), consistent with the acquisition strategy employed in Liu et al. (2024).
@@
Table 2: Prompts used across different stages of LLAMBO and LLAMBO-light in the BBO task.
E.2 Experiment Details for Hyperparameter Tuning Task
@@
RF-4DD
  • max_depth (Maximum depth of a tree): [1,50][-1,50] (integer; 1-1 indicates no limit)
@@
The prompt settings for both LLAMBO and LLAMBO-light in the hyperparameter tuning task follow the same configuration as in the BBO task ($\alpha$ and AF), with the exception of the prompt structure. In particular, the hyperparameter tuning prompts also require both the Description Card and the Data Card to capture the relevant model specifications and historical evaluations.
@@
Appendix F 3D Printing Details
@@
F.1 Qualifying the Stringing Percentage
@@
F.2 Prompts Design
The settings of LLMs are the same as in Appendix E.1. The system prompt is You are an AI assistant that helps me optimize the 3D manufacturing process by controlling parameters. An example of the Data Card is "(Nozzle Temperature, Z Hop Height, Coasting Volume, Retraction Distance, Outer Wall Wipe Distance): (235, 0.3, 0.06, 4, 0.3), Stringing percentage: 12%. We also need a Parameter Description Card to describe the controllable and fixed variables, which is